OpenStudy (anonymous):

Find all the solutions of sin^2 3x- sin^2 x=0 on the interval [0,2pi)

OpenStudy (mertsj):

Are you sure it is sin^2 (3x) ?

OpenStudy (anonymous):

Yup, it's all together.

OpenStudy (mertsj):

Then maybe the second term is also sin^2 (3x) Please check.

OpenStudy (anonymous):

Nope. That's what it says on the paper. I separated the 3x so it wouldn't seem as if it was part of the exponent.

OpenStudy (mertsj):

Then you will have to make a substitution for sin^2 (3x)

OpenStudy (anonymous):

How?O.o

OpenStudy (anonymous):

sin^2 (3x) = sin^2 (x) 3x=x or 3x=-x mod pi

OpenStudy (anonymous):

3x=x or 3x=-x mod pi 2x=0 or 4x=0 mod pi x=n*pi/4, where n is an integer

OpenStudy (anonymous):

x={0, pi/4, pi/2, 3pi/4, pi, 5pi/4, 3pi/2, 7pi/4}

OpenStudy (anonymous):

(pi is the period of sin^2(x); sin^2(-x)=sin^2(x))

OpenStudy (anonymous):

Thank you so much!:) So, what does mod mean?

OpenStudy (anonymous):

mod allows variables to take on multiple values x mod 10, for example, is the set {...x-30,x-20,x-10,x,x+10,x+20,x+30,...} the equations 3x=x or 3x=-x are taken modulus pi since sin^2(x+pi)=sin^2(x) this way, all values of x satisfying that equation sin^2 (3x) = sin^2 (x) will be found